OpenStudy (trojanpoem):

Solid circular disk of radius (0.5 m) rotates around its center of mass under effect of tangential force (F) resulted from pulling a robe coiled around the disk. The angular displacement is given according to the following relation: θ(t)=5t^3+8t^2+4t+10 Given that the torque effecting the disk at t = 2s is 11.4 N.m Find the following: A) Mass of the disk : (Ok) a) 1.2 kg b) 0.15 kg c) 0.3kg d) 0.78 kg B) The force (F) at t = 3s: a)15.9 N b) 31.8 N c) 63.6 N d) 35 N C) Average angular acceleration during t = 4, t =0 D) Length of pulled robe during t = 4, t= 0:

OpenStudy (trojanpoem):

I manged to solve the rest, but got different answers than the aimed. ( in choices, but not correct).

OpenStudy (priyar):

for which q u didn't get the answer?

OpenStudy (trojanpoem):

B, C, D :(

OpenStudy (priyar):

first B

OpenStudy (priyar):

how can u relate torque and force?

OpenStudy (trojanpoem):

M = 1.2 kg

OpenStudy (trojanpoem):

I got the force as a function of time. and let t = 3 s

OpenStudy (trojanpoem):

\[\theta (t) = 5t^{3} + 8 t^{2} + 4t + 10\]

OpenStudy (priyar):

i was about to ask that

OpenStudy (trojanpoem):

My bad.

OpenStudy (priyar):

u got the force as a function of time?

OpenStudy (trojanpoem):

I did.

OpenStudy (priyar):

then what is the problem?

OpenStudy (trojanpoem):

My answer differs from the book.

OpenStudy (trojanpoem):

Write your solution and then we can compare it to mine.

OpenStudy (priyar):

sorry i m not sure..pls check with someone else

OpenStudy (trojanpoem):

@dan815

OpenStudy (priyar):

well..did u differentiate theta to find the angular acceleration and find it 2s?

OpenStudy (trojanpoem):

yeah

OpenStudy (priyar):

then did u divide by r ?

OpenStudy (priyar):

to get the linear acceleration..??

OpenStudy (trojanpoem):

\[\theta (t) = 5t^3 + 8t^2 + 4t +10\] \[\omega (t) = 15t^2 + 16t +4\] \[\omega ' (t) = 30t + 16\] \[F = ma_{t}\] \[a_{t} = \omega r \] \[F = mr \omega = 0.6 (30t + 16) = 18t+ 9.6\] \[F(t) = 18t + 9.6\]

OpenStudy (trojanpoem):

\[F(3) = 18 * 3 + 9.6 = 63.6 N\]

OpenStudy (trojanpoem):

Included in the choices, but not the correct answer.

OpenStudy (priyar):

even i did the same way..so i think ur answer is right

OpenStudy (priyar):

but i took the mass u gave is that correct?

OpenStudy (priyar):

??

OpenStudy (trojanpoem):

Check yourself.

OpenStudy (trojanpoem):

The book has full solution, not just final answer.

OpenStudy (trojanpoem):

I recalculated all his steps and he keep getting 31.8 N which is 0.5 F(3) mine.

OpenStudy (dan815):

umm what ddid u get for the mass?

OpenStudy (trojanpoem):

1.2 kg

OpenStudy (priyar):

ya i checked it..its 1.2kg

OpenStudy (priyar):

where has the book deviated from the above steps?

OpenStudy (trojanpoem):

Book solution for (B): \[\tau = Fr\] \[\tau = I \omega '\] \[\omega ' (3) = 30(3) + 16 = 106 rad/s^2\] \[\tau (3) = 0.15 * 106 = 15.9 N.m\] \[F = \frac{ \tau }{ r } = \frac{ 15.9 }{ 0.5 } = 31.8 N\]

OpenStudy (dan815):

oh ur off by a factor of 2 hmm

OpenStudy (trojanpoem):

or r^2

OpenStudy (trojanpoem):

if it was F = mr^2 omega

OpenStudy (priyar):

hey u have substituted w' in place of w

OpenStudy (trojanpoem):

Typo it must be a = w' r F = w' mr

OpenStudy (priyar):

oh ya..

OpenStudy (priyar):

but does this condition hold true always?

OpenStudy (trojanpoem):

\[s = \theta r \] \[\frac{ d }{ dt }\] \[v = \omega r\] \[\frac{ d }{ dt }\] \[a_{t} = \omega ' r\]

OpenStudy (trojanpoem):

In the proof, only r (the radius) was constant. and it is.

OpenStudy (priyar):

No i am asking like..only in pure rolling a=w'/r is valid right?

OpenStudy (trojanpoem):

Nah, it always hold true. as  S = arc length θ = angular displacement r = raduis S = θr

OpenStudy (trojanpoem):

the one of pure rolling you're talking about is M = a_{r} * N ?

OpenStudy (trojanpoem):

Later in the question, it was used to find D

OpenStudy (priyar):

i didn't get ur notation..a_{r}??

OpenStudy (trojanpoem):

\[M = a_{r} N\]

OpenStudy (priyar):

what does N stand for?

OpenStudy (trojanpoem):

Normal force.

OpenStudy (trojanpoem):

Reaction of the surface upon a circular tire.

OpenStudy (priyar):

ok and M?

OpenStudy (trojanpoem):

The moment rotating the body.

OpenStudy (priyar):

its seems as though both methods are correct...but the answers come out to be different.. if i get any idea i'll tell u

OpenStudy (farcher):

This is a rotational dynamics question with non-uniform angular acceleration. Check my working!

OpenStudy (trojanpoem):

@Farcher , Nice, your steps are totally correct. Can you spot the mistake in my own solution ?

OpenStudy (priyar):

ya @Farcher great job! but why can't we do \[a=\alpha*r\] ?

OpenStudy (priyar):

eventhough the acceleration is non-uniform we are finding its value only at a particular instant..right?

OpenStudy (farcher):

a = alpha r is correct but it is not constant acceleration.

OpenStudy (trojanpoem):

Well, I found the force as a function of time, the solution is above check it.

OpenStudy (farcher):

Then\[\tau = I_c \alpha \Rightarrow Fr = \frac 1 2 m r^2 \alpha \Rightarrow F = \ \frac 12 m a\]

OpenStudy (trojanpoem):

Nice proof, but why isn't the force F = ma as usual ?

OpenStudy (priyar):

i think Farcher is right..here we can't straight away apply F=ma coz the object is under rotational motion..inertia is the rotational equivalent of mass

OpenStudy (farcher):

Because it is not a point mass. It is a disc which is a lot of point masses distributed over a volume. There must actually be another equal and opposite force acting at the axle so that the disc just undergoes rotation and its centre of mas does not move.

OpenStudy (priyar):

Thanks for the explanation..

OpenStudy (trojanpoem):

I now got it , F = ma , resultant of forces = ma ( newton law) i used it as the sole force.

OpenStudy (trojanpoem):

Thanks Farcher.